OpenStudy (anonymous):

A jar contains 14 nickels, 10 dimes, 6 quarters, and 22 pennies. A coin is chosen at random from the jar. What is the probability that the coin chosen is a dime? A. B. C. D.

geerky42 (geerky42):

\(\operatorname{Probability = \dfrac{\operatorname{desired outcomes}}{\operatorname{total outcomes}}}\)

geerky42 (geerky42):

desired outcomes is 10, because there are 10 dimes. total outcomes is 14+10+6+22 = 52

geerky42 (geerky42):

So you have \(\dfrac{10}{52} = \ldots?\)